# Gradient-based Local Next-best-view Planning for Improved Perception of Targeted Plant Nodes

### v1

@@ VI-C Results of occlusion-handling behaviour
TABLE I: Results for the occlusion-handling behaviour. The average performance of 16 experiments is shown. We observed that GradientNBV and SamplingNBV were able to explore the target ROI equally well, but the GradientNBV used ten times less ray-tracing calls and generated four times more efficient trajectories. View 0 was from a predefined global planner.
@@
GradientNBV 10.2 75.2 89.7 92.6 93.8 0 5 10 15 20 0.000 0.151 0.233 0.275 0.308
SamplingNBV 12.9 72.3 90.3 93.7 95.1 0 50 100 150 200 0.000 0.451 0.840 1.251 1.670
Random 7.3 45.1 58.0 63.7 70.4 - - - - - 0.000 0.399 0.851 1.295 1.698
(a) GradientNBV
(b) SamplingNBV
@@
TABLE II: Results for node reconstruction and position estimation. The planners had to detect and improve the perception accuracy of plant nodes within a target ROI. The average performance of $288$ experiments is shown. View 0 was from a predefined global planner.
@@
Trajectory distance (m) $\downarrow$ Recall of occluded node detection $\uparrow$ Standard deviation of 3D node position (m) $\downarrow$
@@
GradientNBV 0.0 32.5 52.8 67.1 78.5 88.6 0.0 0.18 0.28 0.35 0.39 0.42 0.017 0.016 0.014 0.013 0.013 0.012
SamplingNBV 0.0 29.0 54.7 79.1 101.5 123.4 0.0 0.08 0.16 0.21 0.24 0.29 0.017 0.016 0.015 0.014 0.014 0.013
Random 0.0 28.6 54.2 79.4 103.6 127.2 0.0 0.08 0.14 0.16 0.19 0.20 0.017 0.016 0.015 0.015 0.015 0.014
@@ VI-D Results of node reconstruction and position estimation
Figure 3: Example of three consecutive views from GradientNBV. The top row shows result of semantic segmentation and the bottom row shows the rendered information gain (yellow: high gain, blue: low gain). We can observe that the nodes were detected in the third view as the camera moved closer. The gain of the viewpoints within the ROI reduced with each viewpoint.
@@
In this experiment, we tested how well the gradient-based NBV planner handled occlusions in plants. In particular, we analysed if the planner could improve the 3D reconstruction and position estimation of a target node on the plant. We placed a plant in front of the robot with an uncertainty of ±$\pm 10$cm in the y and z axes. An arbitrary node was selected either at the bottom, middle, or top of the plant. An ROI of size $6$cm was defined around this target node with an uncertainty of ±$\pm 3$cm, so that the true position of the node was still within the ROI. This setup mimics the situation in a greenhouse where the 3D position of a node is roughly known, either through direct observation or expectation, and the robot aims to improve its accuracy for the harvesting or de-leafing tasks. Hence, the objective of the view planners was to plan a set of viewpoints that would improve the perception of the target node. Multiple experiments were conducted using $8$ plants with $3$ target nodes and $12$ rotations along the z-axis per plant, leading to a total of $288$ experiments per planner. The planning was terminated at the end of $5$ viewpoints.
@@
In Table II, we quantitatively analysed the planner performances. We observed that the GradientNBV and SamplingNBV planners performed equally well in ROI coverage and F1-score, which indicated that both planners were able to handle occlusion and improve the 3D reconstruction of the target nodes. However, GradientNBV was ten times more efficient in terms of computational time and $28\%$ more efficient in terms the camera trajectory. Regarding the detection of occluded nodes across all experiments, the recall of GradientNBV was $13\%$ more than SamplingNBV and $22\%$ more than Random, indicating that the GradientNBV was effective in detecting the subset of nodes that were undetected in the starting view. The estimated standard deviation of the 3D node positions were low and reduced with more viewpoints for all the planners. Fig. 3 visualises the node detection and the rendered information gain for views generated by GradientNBV.
VII Conclusion
@@
In this paper, we presented a novel gradient-based NBV planning method using differential ray sampling for improving the perception of target objects in occluded scenarios. Compared to previous works on NBV planning, our approach removes the need to sample and evaluate multiple candidate viewpoints. We performed simulation experiments with 3D mesh models of tomato plants to evaluate the performance of our planner on the tasks of 3D reconstruction and position estimation of nodes. Our gradient-based NBV planner was able to explore the target region, reconstruct the nodes, and estimate their 3D position as accurately as a sampling-based NBV planner, while taking ten times less computation and generating $28\%$ more efficient trajectories. Our results clearly show the advantage of using the gradient-based NBV planner for local viewpoint planning in occluded scenarios. Our approach can help robots to efficiently perceive a target plant node for grasping or cutting, and contribute to improved harvesting and de-leafing in greenhouses.

### v2

@@ VI-C Results of occlusion-handling behaviour
TABLE I: Results for the occlusion-handling behaviour. The average performance across 16 experiments is shown. We observed that GradientNBV and SamplingNBV were able to explore the target ROI equally well, but the GradientNBV used ten times less ray-tracing calls and generated four times more efficient trajectories. View 0 was predefined.
@@
GradientNBV 10.2 75.2 89.7 92.6 93.8 0 5 10 15 20 0.00 0.15 0.23 0.28 0.31
SamplingNBV 12.9 72.3 90.3 93.7 95.1 0 50 100 150 200 0.00 0.45 0.84 1.25 1.67
Random 7.3 45.1 58.0 63.7 70.4 - - - - - 0.00 0.40 0.85 1.29 1.70
@@
TABLE II: Results for node reconstruction and position estimation of plant nodes within a target ROI. We observed that the GradientNBV and SamplingNBV explored the target ROI equally well, but the GradientNBV used ten times less ray-tracing calls and generated $28\%$ more efficient trajectories. The average performance across $288$ experiments is shown. View 0 was predefined.
@@
Trajectory distance (m) $\downarrow$ Recall of occluded node detection (%) $\uparrow$ $\sigma$ of 3D node position (m×$\times 10^{-2}$) $\downarrow$
@@
GradientNBV 0.0 0.11 0.18 0.23 0.27 0.31 0.0 18.2 28.5 35.2 39.4 41.8 1.7 1.6 1.4 1.3 1.3 1.2
SamplingNBV 0.0 0.10 0.19 0.28 0.35 0.43 0.0 7.5 16.2 20.6 24.4 29.4 1.7 1.6 1.5 1.4 1.4 1.3
Random 0.0 0.10 0.19 0.28 0.36 0.44 0.0 8.4 14.4 16.3 18.7 20.5 1.7 1.6 1.5 1.5 1.5 1.4
@@ VI-D Results of node reconstruction and position estimation
Figure 3: Example of three consecutive views from GradientNBV. The top row shows the segmented images and the bottom row shows the viewpoint utility rendered from the current view (yellow: high, blue: low). We can observe that the nodes were detected in View 3 as the camera moved closer. The utility of the viewpoints within the ROI reduced with each viewpoint.
@@
In this experiment, we tested how well the gradient-based NBV planner handled occlusions in plants. In particular, we analysed if the planner could improve the 3D reconstruction and position estimation of a target node on the plant. We placed a plant in front of the robot with an uncertainty of ±$\pm 10$cm in the y and z axes. An arbitrary node was selected either at the bottom, middle, or top of the plant. An ROI of size $6$cm was defined around this target node with an uncertainty of ±$\pm 3$cm, so that the true position of the node was still within the ROI. $p^{t}$ was initialised with the target node’s position and $p^{c}$ was predefined close to it. This setup mimicked the situation in a greenhouse where the 3D position of a node was roughly known, either through direct observation or expectation, and the robot aimed to improve its accuracy for harvesting or de-leafing. Hence, the objective was to plan a set of viewpoints that would improve the perception of the target node. Multiple experiments were conducted using $8$ plants with $3$ target nodes and $12$ rotations along the z-axis per plant, leading to a total of $288$ trials per planner. In $165$ trials, the nodes were occluded and undetected in the starting view. The planning was terminated at the end of $5$ viewpoints.
@@
In Table II, we quantitatively analysed the planner performances. We observed that the GradientNBV and SamplingNBV planners performed equally well in ROI coverage and F1-score, which indicated that both planners were able to handle occlusion and improve the 3D reconstruction of the target nodes. However, GradientNBV was ten times more computationally efficient and $28\%$ more efficient in terms of the camera trajectory. Regarding the detection of occluded nodes, the recall of GradientNBV was $13\%$ more than SamplingNBV and $22\%$ more than Random, indicating that the GradientNBV was more effective in detecting the nodes that were undetected in the starting view. The estimated standard deviation of the 3D node positions were low and reduced with more viewpoints for all the planners. Fig. 3 visualises the node detection and the rendered utility for views generated by GradientNBV.
VII Conclusion and Future work
@@
In this paper, we presented a novel gradient-based NBV planning method using differentiable ray sampling for improving the perception of target objects in occluded scenarios. Compared to previous works on NBV planning, our approach removes the need to sample and evaluate multiple candidate viewpoints. We performed simulation experiments with 3D mesh models of tomato plants to evaluate the performance of our planner on the tasks of 3D reconstruction and position estimation of nodes. Our gradient-based NBV planner was able to explore the target region, reconstruct the nodes, and estimate their 3D position as accurately as a sampling-based NBV planner, while taking ten times less computation and generating $28\%$ more efficient trajectories. Our results clearly show the advantage of using the gradient-based NBV planner for local viewpoint planning in occluded scenarios. Our approach can help robots to efficiently perceive a target plant node for grasping or cutting, and contribute to improved harvesting and de-leafing in greenhouses. In future work, we can extend the method for global planning in larger scenarios.